# BrainMorph: A Foundational Keypoint Model for Robust and Flexible Brain MRI Registration

### v1

@@ Abstract
{IEEEkeywords}
Image registration, Multi-modal, Keypoint detection, Foundation model, Brain MRI
1 Introduction
@@
Figure 1: The framework of BrainMorph. Fixed and moving 3D brain images are passed through the same keypoint detection network, which predicts $N$ keypoints useful for registration. The transformation parameters are then computed as a function of the keypoints, which are in turn used to resample the moving image. Keypoint colors denote depth (see Fig. 2).
@@ 3 BrainMorph
Our tool, BrainMorph, is based on the previously published KeyMorph framework and we refer the reader to prior papers for more details [15, 16]. Let $(\bm{x}_{m},\bm{x}_{f})$ be a moving (source) and fixed (target) image [^1] Although we consider 3D volumes in this work, BrainMorph is agnostic to the number of dimensions. The terms “image” and “volume” are used interchangeably. pair, possibly of different contrasts or modalities. Additionally, we denote by $\mathcal{T}_{\theta}$ a parametric coordinate transformation with parameters $\theta$, such as those discussed in Appendix .1. The goal is to find the optimal transformation $\mathcal{T}_{\theta^{*}}$ such that the registered image $\bm{x}_{r}=\bm{x}_{m}\circ\mathcal{T}_{\theta^{*}}$ aligns with some fixed image $\bm{x}_{f}$, where $\circ$ denotes the spatial transformation of an image.
@@ 3.4 Groupwise Registration
Figure 2: Moving, fixed, and aligned images in axial, sagittal, and coronal mid-slices. Keypoints extracted by BrainMorph are overlaid. For Aligned, aligned keypoints are dots and fixed keypoints are crosses. Color of the keypoint corresponds to depth. Note that keypoint weights are not visualized.
@@ 4.3 Model Details
ANTs, Rigid 101.38$\pm$2.33 - 143.38$\pm$2.33 -
ANTs, Affine 110.45$\pm$2.94 - 142.45$\pm$2.94 -
ANTs, Syn 216.03$\pm$3.14 - 248.03$\pm$3.14 -
Table 4: Average computation time in seconds for pairwise registration across different models. We separate between times when performing registration with pre-processed data and raw data. For ANTs, pre-processing requires skull-stripping with HD-BET, which adds an additional 32 seconds [68]. For SynthMorph, pre-processing requires an initial robust affine registration into a reference space. KeyMorph timings are based on KeyMorph-S.
@@ 5.3 Metrics
Figure 3: Dice performance on pairwise registration. Higher is better. Unimodal/multimodal, skull-stripped/non-skull-stripped.
Figure 4: HD performance on pairwise registration. Lower is better. Unimodal/multimodal, skull-stripped/non-skull-stripped.
@@ 5.4 Baselines
Figure 5: Pairwise registration results for BrainMorph and selected baselines. In first row, a 90 degree rotation is applied. In second row, a 135 degree rotation is applied.
@@ 6.1.2 Groupwise Registration
Figure 6: Groupwise registration with group size of 4. For ITK-Elastix, B-spline registration is used. For BrainMorph, TPS with $\lambda=0$ is used. The last column shows the average brain in the optimized template space.
Figure 7: Groupwise registration using BrainMorph and ITK-Elastix. For BrainMorph, the transformation used is TPS with $\lambda=0$. For ITK-Elastix, the transformation is nonlinear b-spline. We observe significant improvement in groupwise registration capabilities of BrainMorph, at a time reduction of nearly 4 orders of magnitude.
@@ 6.1.3 Longitudinal Registration
Figure 8: a) Registration performance on longitudinal registration. Rigid transformations for all models. b) Registration performance on subjects with lesions. Affine transformations for all models, except for SynthMorph which only supports dense.
@@ 6.2.2 Number of Keypoints
Figure 9: Registration performance of variants of BrainMorph and varying the number of keypoints.

### v2

@@ Abstract
keywords:
@@
1 Introduction
Fig. 1: The framework of BrainMorph. Fixed and moving 3D brain images are passed through the same keypoint detection network, which predicts $N$ keypoints useful for registration. The transformation parameters are then computed as a function of the keypoints, which are in turn used to resample the moving image. Keypoint colors denote depth (see Fig. 2).
Fig. 2: Moving, fixed, and aligned images in axial, sagittal, and coronal mid-slices. Keypoints extracted by BrainMorph are overlaid. For Aligned, aligned keypoints are dots and fixed keypoints are crosses. Color of the keypoint corresponds to depth. Note that keypoint weights are not visualized.
@@ 2 Background
ANTs, Rigid 101.38$\pm$2.33 - 143.38$\pm$2.33 -
ANTs, Affine 110.45$\pm$2.94 - 142.45$\pm$2.94 -
ANTs, Syn 216.03$\pm$3.14 - 248.03$\pm$3.14 -
Table 2: Average computation time in seconds for pairwise registration across different models. We separate between times when performing registration with pre-processed data and raw data. For ANTs, pre-processing requires skull-stripping with HD-BET, which adds an additional 32 seconds [41]. For SynthMorph, pre-processing requires an initial robust affine registration into a reference space. KeyMorph timings are based on KeyMorph-L.
@@ 3 BrainMorph
Our tool, BrainMorph, is based on the previously published KeyMorph framework and we refer the reader to prior papers for more details [98, 92]. Let $(\bm{x}_{m},\bm{x}_{f})$ be a moving (source) and fixed (target) image [^1] Although we consider 3D volumes in this work, BrainMorph is agnostic to the number of dimensions. The terms “image” and “volume” are used interchangeably. pair, possibly of different contrasts or modalities. Additionally, we denote by $\mathcal{T}_{\theta}$ a parametric coordinate transformation with parameters $\theta$, such as those discussed in A. The goal is to find the optimal transformation $\mathcal{T}_{\theta^{*}}$ such that the registered image $\bm{x}_{r}=\bm{x}_{m}\circ\mathcal{T}_{\theta^{*}}$ aligns with some fixed image $\bm{x}_{f}$, where $\circ$ denotes the spatial transformation of an image.
@@ 4.4 Self-supervised Pretraining
Fig. 3: Dice performance on pairwise registration. Higher is better. Unimodal/multimodal, skull-stripped/non-skull-stripped.
Fig. 4: HD performance on pairwise registration. Lower is better. Unimodal/multimodal, skull-stripped/non-skull-stripped.
@@ 5 Experimental Setup
Fig. 5: Pairwise registration results for BrainMorph and selected baselines. In first row, a 90 degree rotation is applied. In second row, a 135 degree rotation is applied.
Fig. 6: Groupwise registration with group size of 4. For ITK-Elastix, B-spline registration is used. For BrainMorph, TPS with $\lambda=0$ is used. The last column shows the average brain in the optimized template space.
@@ 5.1 Evaluation Datasets
Fig. 7: Groupwise registration using BrainMorph and ITK-Elastix. For BrainMorph, the transformation used is TPS with $\lambda=0$. For ITK-Elastix, the transformation is nonlinear b-spline. We observe significant improvement in groupwise registration capabilities of BrainMorph, at a time reduction of nearly 4 orders of magnitude.
@@ 5.3 Metrics
Fig. 8: a) Registration performance on longitudinal registration. Rigid transformations for all models. b) Registration performance on subjects with lesions. Affine transformations for all models, except for SynthMorph which only supports dense.
@@ 5.4 Baselines
  • 1.
@@ 6.2.2 Number of Keypoints
Fig. 9: Registration performance of variants of BrainMorph and varying the number of keypoints.
@@ 8 Conclusion
9 Acknowledgments
We thank Benjamin Billot and Juan Eugenio Iglesias for helpful discussions. Funding for this project was in part provided by the NIH grants R01AG053949, R01AG064027 and R01AG070988, and the NSF CAREER 1748377 grant.
Appendix A Differentiable, Closed-Form Coordinate Transformations
Notation: In the following sections, column vectors are lower-case bolded and matrices are upper-case bolded. $D$-dimensional coordinates are represented as column vectors, i.e. $\bm{p}\in\mathbb{R}^{D}$. $D$ is typically 2 or 3. $\tilde{\bm{p}}$ denotes $\bm{p}$ in homogeneous coordinates, i.e. $\tilde{\bm{p}}=[\bm{p},1]^{T}$. Superscripts in parentheses $\bm{p}^{(i)}$ index over separate instances of $\bm{p}$ (e.g. in a dataset), whereas subscripts $\bm{p}_{i}$ denotes the $i$’th element of $\bm{p}$.
We summarize three parametric transformation families that can be derived in closed-form, from corresponding keypoint pairs. Suppose we have a set of $N$ corresponding keypoint pairs $\{(\bm{p}^{(i)},\bm{q}^{(i)})\}_{i=1}^{N}$, where $\bm{p}^{(i)},\bm{q}^{(i)}\in\mathbb{R}^{D}$ and $N>D$. For convenience, let $\bm{P}:=\begin{bmatrix}\bm{p}^{(1)}&...&\bm{p}^{(N)}\end{bmatrix}\in\mathbb{R}^{D\times N}$, and similarly for $\tilde{\bm{P}}$ and $\bm{Q}$. Define $\mathcal{T}_{\theta}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ as a family of coordinate transformations, where $\theta\in\Theta$ are parameters of the transformation. For all transformation families, we also consider weighted versions, where we have weights for each correspondence $\{w_{i}\}_{i=1}^{N}$. For convenience, let $\bm{W}=\text{diag}(w_{1},...,w_{N})$.
A.1 Rigid
Rigid transformations apply a rotation $\bm{R}\in\mathbb{R}^{D\times D}$ and a translation $\bm{t}\in\mathbb{R}^{D\times 1}$ to a coordinate:
$\mathcal{T}_{\theta}(\bm{p})=\bm{R}\bm{p}+\bm{t},$ (4)
where the parameter set is the elements of the matrix and vector, $\theta=\{\bm{R},\bm{t}\}$.
The optimal translation is estimated by subtracting the weighted centroids of the moving and fixed point clouds:
$\bm{t}^{*}=\bar{\bm{p}}-\bar{\bm{q}},$ (5)
where $\bar{\bm{p}}=\sum_{i}\bm{p}^{(i)}$ for non-weighted and $\bar{\bm{p}}=\sum_{i}w_{i}\bm{p}^{(i)}$ for weighted, and similarly for $\bar{\bm{q}}$.
The optimal rotation is well-studied and is known as the orthogonal Procrustes problem [88]. First, denote by $\bar{\bm{P}}$ the centered version of $\bm{P}$ where each column is subtracted by the centroid $\bar{\bm{p}}$, and similarly for $\bar{\bm{Q}}$. Next, compute the SVD of the weighted cross-correlation matrix $\text{SVD}(\bm{\Sigma})=\text{SVD}(\bar{\bm{P}}^{T}\bar{\bm{Q}})=\bm{U}\bm{\Lambda}\bm{V}^{T}$. For weighted, $\Sigma=\bar{\bm{P}}^{T}\bm{W}\bar{\bm{Q}}$. Then, $\bm{R}^{*}=\bm{V}\bm{U}^{T}$.
A.2 Affine
Affine transformations are represented as a matrix multiplication of $\bm{A}\in\mathbb{R}^{D\times(D+1)}$ with a coordinate in homogeneous form:
$\mathcal{T}_{\theta}(\bm{p})=\bm{A}\tilde{\bm{p}},$ (6)
where the parameter set is the elements of the matrix, $\theta=\{\bm{A}\}$.
Given $N$ corresponding keypoint pairs, there exists a differentiable, closed-form expression for an affine transformation that aligns the keypoints:
$\theta^{*}(\bm{P},\bm{Q})$ $:=\operatorname*{arg\,min}_{\theta}\sum_{i=1}^{N}\left(\bm{A}\tilde{\bm{p}}^{(i)}-\bm{q}^{(i)}\right)^{2}$ (7)
$=\bm{Q}\tilde{\bm{P}}^{T}(\tilde{\bm{P}}\tilde{\bm{P}}^{T})^{-1}.$ (8)
To derive this solution, rewrite the objective in matrix form:
$\mathcal{L}$ $=\sum_{i=1}^{N}\left(\bm{A}\tilde{\bm{p}}^{(i)}-\bm{q}^{(i)}\right)^{2}$
$=\left\|\bm{A}\tilde{\bm{P}}-\bm{Q}\right\|_{F},$
where $\left\lVert\cdot\right\rVert_{F}$ denotes the Frobenius norm. Taking the derivative with respect to $\bm{A}$ and setting the result to zero, we obtain:
$\frac{\partial\mathcal{L}}{\partial\bm{A}}$ $=(\bm{A}\tilde{\bm{P}}-\bm{Q})\tilde{\bm{P}}^{T}=\bm{0}$
$\implies\bm{A}\tilde{\bm{P}}\tilde{\bm{P}}^{T}=\bm{Q}\tilde{\bm{P}}^{T}$
$\implies\bm{A}=\bm{Q}\tilde{\bm{P}}^{T}(\tilde{\bm{P}}\tilde{\bm{P}}^{T})^{-1}.$
The extension to incorporate weightings for the correspondences is straightforward:
$\theta^{*}(\bm{P},\bm{Q},\bm{W})$ $:=\operatorname*{arg\,min}_{\theta}\sum_{i=1}^{N}w_{i}\left(\bm{A}\tilde{\bm{p}}^{(i)}-\bm{q}^{(i)}\right)^{2}$ (9)
$=\bm{Q}\bm{W}\tilde{\bm{P}}^{T}(\tilde{\bm{P}}\bm{W}\tilde{\bm{P}}^{T})^{-1}.$ (10)
Note that solving for the affine transformation is the least-squares solution to an overdetermined system, and thus in practice the points will not be exactly matched due to the restrictive nature of the affine transformation. This restrictiveness may be alleviated or removed by choosing a transformation family with additional degrees of freedom, as we detail next.
A.3 Thin-Plate Spline
The application of the thin-plate spline (TPS) interpolant to modeling coordinate transformations yields a parametric, non-rigid deformation model which admits a closed-form expression for the solution that interpolates a set of corresponding keypoints [11, 19, 73]. This provides additional degrees of freedom over the affine family of transformations, while also subsuming it as a special case.
For the $d$’th dimension, the TPS interpolant $\mathcal{T}_{\theta_{d}}:\mathbb{R}^{D}\rightarrow\mathbb{R}$ takes the following form:
$\mathcal{T}_{\theta_{d}}(\bm{p})=(\bm{a}_{d})^{T}\tilde{\bm{p}}+\sum_{i=1}^{N}v_{i,d}U\left(\left\lVert\bm{p}^{(i)}-\bm{p}\right\rVert_{2}\right),$ (11)
where $\bm{a}_{d}\in\mathbb{R}^{D+1}$ and $\{v_{i,d}\}$ constitute the transformation parameters $\theta_{d}$ and $U(r)=r^{2}\ln(r)$. We define $\bm{A}\in\mathbb{R}^{(D+1)\times D}$ and $\bm{V}\in\mathbb{R}^{N\times D}$ as the collection of all the parameters for $d=1,...,D$. Then, the parameter set is $\theta=\{\bm{A},\bm{V}\}$.
This form of $\mathcal{T}$ minimizes the bending energy:
[…]
where $\left\lVert\cdot\right\rVert_{F}$ is the Frobenius norm and $\nabla^{2}\mathcal{T}$ is the matrix of second-order partial derivatives of $\mathcal{T}$. For each $\theta_{d}$, we impose interpolation conditions $\mathcal{T}_{\theta_{d}}(\bm{p}^{(i)})=\bm{q}_{d}^{(i)}$ and enforce $\mathcal{T}$ to have square-integrable second derivatives:
$\sum_{i=1}^{N}v_{i,d}=0\ \ \text{and}\ \ \sum_{i=1}^{N}v_{i,d}\bm{p}_{d}=0\ \ \forall d\in\{1,...,D\}.$ (13)
Given these conditions, the following system of linear equations solves for $\theta$:
$\bm{\Psi}\theta:=\begin{bmatrix}\bm{K}&\bm{L}\\ \bm{L}^{T}&\bm{O}\end{bmatrix}\begin{bmatrix}\bm{V}\\ \bm{A}\end{bmatrix}=\begin{bmatrix}\bm{Q}^{T}\\ \bm{O}\end{bmatrix}:=\bm{Z}.$ (14)
Here, $\bm{K}\in\mathbb{R}^{N\times N}$ where $\bm{K}_{ij}=U\left(\left\lVert\bm{p}^{(i)}-\bm{p}^{(j)}\right\rVert_{2}\right)$, $\bm{L}\in\mathbb{R}^{N\times(D+1)}$ where the $i$’th row is $(\tilde{\bm{p}}^{(i)})^{T}$, and $\bm{O}$ is a matrix of zeros with appropriate dimensions. Thus,
$\theta^{*}(\bm{P},\bm{Q}):=\bm{\Psi}^{-1}\bm{Z}.$ (15)
Solving for $\theta^{*}$ is a differentiable operation.
The interpolation conditions can be relaxed (e.g. under the presence of noise) by introducing a regularization term:
$\operatorname*{arg\,min}_{\theta_{d}}\sum_{i=1}^{N}\left(\mathcal{T}_{\theta_{d}}\left(\bm{p}^{(i)}\right)-\bm{q}_{d}^{(i)}\right)^{2}+\lambda I_{\mathcal{T}}$ (16)
where $\lambda>0$ is a hyperparameter that controls the strength of regularization. As $\lambda$ approaches $\infty$, the optimal $\mathcal{T}$ approaches the affine case (i.e. zero bending energy). This formulation can be solved exactly by replacing $\bm{K}$ with $\bm{K}+\lambda\bm{I}$ in Eq. (14). Importantly, $\theta$ and the optimal $\theta^{*}(\bm{P},\bm{Q})$ exhibits a dependence on $\lambda$. Finally, weights can be incorporated by replacing $\bm{K}$ with $\bm{K}+\lambda\bm{W}^{-1}$ [73].
Appendix B List of Datasets
A full list of datasets used for training and evaluation is shown in Table 5.
Dataset Description Train/Eval No. Subjects No. Modalities Modalities Info No. Images Longitudinal? Lesion?
BraTS-SSA-2023 African Gliloms Train 60 4 T1 T1gd T2 FLAIR 240
BraTS-MEN-2023 Meningiomas Train 1000 4 T1 T1gd T2 FLAIR 4000
BraTS-MET-2023 Metastases Train 165 4 T1 T1gd T2 FLAIR 660
BraTS-MET-NYU-2023 Metastases Train 164 4 T1 T1gd T2 FLAIR 656
BraTS-PED-2023 Pediatric Glioblastomas Train 99 4 T1 T1gd T2 FLAIR 396
BraTS-MET-UCSF-2023 Metastases Train 324 4 T1 T1gd T2 FLAIR 1296
BraTS-2016 Tumor Train 1104 1 T1 1104
UCSF-BMSR Metastases Train 459 4 FLAIR T1post T1pre subtraction 1836
MSSEG2 Multiple Sclerosis Train 40 1 FLAIR 80
Shifts-challenge-part1MSSEG Multiple Sclerosis Train 15 5 FLAIR T1 T2 GADO DP 75
Shifts-challenge-part2Ljubljana Multiple Sclerosis Train 25 4 FLAIR T1 T2 T1ce 100
Shifts-challenge-part2Best Multiple Sclerosis Train 21 4 FLAIR T1 T2 PD 84
Brain Development Adult and Neonatal Brain Atlases Train 33 1 T1 33
EPISURG Epileptic Train 133 1 T1 133
WMH White Matter Hyperintensity Train 120 1 T1 120
ISLES-2022 Ischemic Stroke Lesion Train 250 2 DWI ADC 500
MedicalDecathlon-BrainTumour Tumor Train 484 4 FLAIR T1 T1gd T2 1936
UCSF-ALPTDG Post Treatment Gliomas Train 298 5 FLAIR T1 T1ce T2 T1csub 1490
StanfordMETShare Metastases Train 105 4 T1 FLAIR T1gd T1pre 420
PPMI Parkinson’s Disease Train 1382 1 T1 1382
ADNI Alzheimer’s Disease Train 2027 3 T1 T2 FLAIR 18645
PAC 2019 Normal Aging Train 2640 1 T1 2640
AIBL Alzheimer’s Disease Train 708 1 T1 708
OASIS1 Normal Aging and Dementia Train 348 1 T1 1350
OASIS3 Normal Aging and Alzheimer’s Disease Train 2270 6 T1 T2 T2star FLASH FLAIR angio 10299
IXI Normal Eval 151 3 T1 T2 PD 453
OASIS2 Dementia Eval 298 1 T1 1095
BraTS-GLI-2023 (test) Glioblastomas Eval 876 4 T1 T1gd T2 FLAIR 3504
MS-60 Multiple sclerosis Eval 60 3 T1 T2 FLAIR 180
Total 15659 13 (unique) 55415 ($\times$ 2)
Table 5: Dataset details. Total images is 55415 (possibly non-unique due to overlap). $\times$ 2 = 110836 to include with and without skullstripping. All citations in top-bottom row order: [42], [42][42][42][42][76][62][76][16][56][76][76][28][69][46][34][1][75][29][59][70][27][24][58][47]https://brain-development.org/ixi-dataset/[57][42][65].